Chuck, a single taxpayer,earns $86,000 in taxable income and $24750 in interest from an investment in City of Heflin bonds. (Use the U.S. tax rate schedule) Required: a. If Chuck earns an additional $41,000 of taxable income, what is his marginal tax rate on this income? b. What is his marginal rate if, instead, he had $41,000 of additional deductions? (For all requirements, do not round intermediate calculations. Round your answers to 2 decimal places.) a. Marginal tax rate b Marginal tax rate 1%
Answer #1

Interest from the investment in the city of heflin bonds (municipal bonds) is tax free

Chuck is a single tax payer . taxable income = $86,000
Therefore, tax payable = 5,183.75 + (86,000 - 37,650) * 25%
= 5,183.75 + 12,087.5
= $ 17,271.25
a) additional 41,000 of taxable income is earned
total taxable income = 86,000 + 41,000 = $ 127,000
Marginal tax rate on the income of 41,000 = 11,325.50 / 41,000 = 27.6232 %
tax on income of 53,000 = (91,150 - 86,000) * 25% + (41,000 - 5,150) * 28 %
= (5150) * 25% + (35850) * 28 %
= 1,287.50 + 10,038
= 11,325.50
Tax on income = 18558.75 + (127,000 - 91,150) * 28%
= 18558.75 + 10,038
= 28596.75
B)
taxable income after deductions of 41,000 = 86,000 - 41,000 = $ 45,000
= 5,183.75 + (45,000 - 37,650) * 25%
= 7,021.25 7350
1837.5 7021.25
marginal tax rate = 7021.25 /45,000 = 15.60 %